# Machine-generated text detection prevents language model collapse

### v4

@@ 4 The impact of decoding strategies on model collapse
We carry out recursive training as described in Section 3.2 on the open-ended text generation task by fine-tuning LLMs on the WikiText-2 dataset Merity et al. (2016) with GPT-2 Radford et al. (2019) and SmolLM2 Allal et al. (2025). The Wikipedia articles are segmented into non-overlapping chunks of $512$ tokens, where the first $256$ are used as the context ($\mathbf{x}$), and the remaining $256$ as the continuation ($\hat{\mathbf{x}}$). We conduct full fine-tuning for $1$ epoch and, to avoid information leakage between generation and training, define cross-entropy loss only on the generated sequence of each sample Dohmatob et al. (2024b). Additional details can be found in Appendix B.3. As described in section 3.2, we recursively train language models on the previous generation’s output. We evaluate a range of decoding strategies to assess the effect on model collapse: greedy decoding ($\tau=0$), $5$-way beam search, pure sampling ($\tau=1$), temperature ($\tau=0.9$), top-$k$ (with $k=50$), and nucleus sampling ($\eta=0.95$). The hyperparameter settings for these decoding strategies were based on recommendations from prior work Holtzman et al. (2020); Shumailov et al. (2024); Garces Arias et al. (2025).
@@
Figure 1 depicts the perplexity and evaluation accuracy on the WikiText-2 test set for every model generation. Additionally, we obtain scores for the qualitative metrics using the outputs generated by the model (i.e. $\{\hat{\mathbf{x}}\}_{s=1}^{n}$ of $\mathcal{D}_{\text{S}}^{i}$ in Algorithm 1), and enumerate them in Table 1 for generations 00 and $9$. We first notice that deterministic decoding results in the most significant model collapse. While stochastic sampling methods appear to follow a linear relationship in model degradation across generations, the rate of collapse is accelerated with greedy decoding and beam search which then starts to plateau in later generations. At generation 00, deterministic strategies yield significantly less fluent and more repetitive text, with MAUVE scores less than $5\%$ and diversity scores less than $20\%$. The gap in generation quality between deterministic and stochastic strategies in the open-ended text generation task has been explored in related literature Holtzman et al. (2020); Pillutla et al. (2021). Here we see how it leads to exacerbated model collapse. While deterministic methods are rarely used in open-ended generation, we included them for completeness and due to the choice of beam-search for the experiments of Shumailov et al. (2024), but exclude them in subsequent experiments due to their unrealistic collapse.
@@
Pure sampling approaches appear to have affected the language models in contrasting ways. For GPT-2, sampling directly from the probability distribution leads to diverse and fluent text in generation 00, but training recursively on these outputs yields the worst test perplexity among all stochastic methods, and generated text that is significantly dissimilar to human outputs (MAUVE of $7.18$). With SmolLM2, on the other hand, pure sampling led to the smallest drop in model task performance and the closest similarity to the human reference on average across all metrics.
Importantly, stochastic decoding with temperature leads to the most repetitive outputs after recursive training. We observed a ${}70\%$ decrease in diversity for both models and the farthest distribution semantically from the human reference for SmolLM2, showing significant model collapse. Performance with top-$k$ sampling, on the other hand, was consistent across models, with the smallest drop in diversity and Self-BLEU, the closest resulting text to the human reference, and a smaller drop in test performance compared to nucleus sampling.
@@
In our subsequent experiments on preventing model collapse, we seek to validate that our method can work in the most extreme scenario while also facilitating direct comparison between models. For this reason, we evaluate the models using the worst-performing stochastic decoding method (pure sampling for GPT-2 and temperature sampling for SmolLM2) and top-$k$ decoding due to its consistent performance across models.
@@ 5 Preventing model collapse
So far we have carried out recursive training in a setting where models are trained exclusively on the outputs of the previous generation without implicitly including any human-generated samples. We now turn our focus to the partially synthetic setting, a more realistic scenario where human data make up a portion of the training dataset and the synthetic data is a mix of the samples produced across generations. The training dataset for generation $i$, $\mathcal{D}^{i}$, consists of the aggregation of $3$ samples:
@@
We explore the following settings: (i) fully synthetic $\left(\alpha\!=\!0,\beta\!=\!1,\gamma\!=\!0\right)$, where training data consists entirely of synthetic samples from the previous generation, (ii) partially synthetic $\left(\alpha\!>\!0,\beta\!=\!1,\gamma\!=\!0\right)$, where the same proportion of human data is added to the training data at every generation, and (iii) partially synthetic with cross-generational data $\left(\alpha\!>\!0,\beta\!>\!0,\gamma\!>\!0\right)$ which is the same as above, but with the synthetic data accumulated across generations. We evaluate our method in the partially synthetic setting and vary the mixing coefficients $\alpha$, $\beta$, and $\gamma$, however, our method does not assume access to the values of the mixing coefficients and hence the data distribution. To prevent model collapse when the origin of each training sample is unknown, we train a machine-generated detector that estimates the likelihood of text origin (section 5.1). We then use this information to conduct weighted sampling (section 5.2) that ultimately mitigates model collapse.
@@
5.1 Machine-generated text detection performance
We trained and evaluated RoBERTa, DeBERTav3 and ModernBERT on the MAGE dataset Li et al. (2024), a machine-generated text detection benchmark based on documents from $10$ domains which have been used to generate text from $27$ LLMs. We adopt the preset training / validation / test splits ($80\%$/$10\%$/$10\%$). We also test on the more demanding out-of-distribution test set that contains human-curated text from $4$ unseen domains and machine-generated samples by an unseen LLM (GPT-4). Each model was fine-tuned for 5 epochs using a binary cross-entropy loss with early stopping. More details on the model training can be found in Appendix A. Performance is enumerated in Table 2. ModernBERT yielded the best classification performance on both aforementioned test sets with an AUC of .986.986.986.986 and .943.943.943.943, respectively. This is comparable to the top-performing model evaluated by Li et al. (2024), Longformer, which achieved an in-distribution and out-of-distribution AUC of .99.99.99.99 and .94.94.94.94, respectively. We release the fine-tuned ModernBERT model publicly [^3] https://huggingface.co/GeorgeDrayson/modernbert-ai-detection.
@@ 5.2 Informed sampling of training data
Our goal is to sample a new training dataset at each iteration of recursive training which does not result in model collapse. We consider that a language model has collapsed if the addition of synthetic samples in the training data leads to degraded model performance compared to training exclusively on the human samples. Given text inputs $\mathbf{x}_{i}$, $i\!\in\!\{1,\dots,n\}$, with an unknown distribution of human and synthetic samples, the machine-generated text detector yields predictions $q\!\left(\mathbf{x}_{i}\right)$, $\forall i$.
Our approach deploys weighted sampling with replacement, using the detector’s prediction probabilities as weights. As the detector has been trained on an unbalanced dataset ($29\%$ human samples), the predictions are biased towards attributing text as machine-generated, reflected in the optimal classification threshold of $0.8674$ (00/$1$: human/machine-generated). To ameliorate this, we apply a bias term $b\!\geq\!1$ (see Appendix C.1) to the probabilities, followed by normalising the weights using
@@
where $w_{i}\!\in\![0,1]$ denotes the weight for sample $\mathbf{x}_{i}$, $\forall i$. From the $n$ weighted training samples, we draw $k\!\times\!n$ samples with replacement. We have set $k$ to $1.5$ to allow for a $50\%$ upsampling of the training data. To avoid oversampling of specific samples we also set the maximum number of selections for an individual sample to $10$. In this way, we obtain a revised set of samples that we use in our recursive training regime.
@@
sampling .5.5.5.5, $1$, 00 ${\color[rgb]{0,0,1}-7.41\%}$ ${\color[rgb]{0,0,1}+1.50\%}$ ${\color[rgb]{1,0,0}-1.30\%}$ ${\color[rgb]{1,0,0}+36.71\%}$ ${\color[rgb]{0,0,1}+74.06\%}$ ${\color[rgb]{0,0,1}+50.23\%}$
@@
SmolLM2 top-$k$5.5.5.5, $1$, 00 ${\color[rgb]{0,0,1}-3.56\%}$ ${\color[rgb]{0,0,1}+0.40\%}$ ${\color[rgb]{0,0,1}+10.51\%}$ ${\color[rgb]{0,0,1}-5.42\%}$ ${\color[rgb]{0,0,1}+2.29\%}$ ${\color[rgb]{1,0,0}-2.50\%}$
.5.5.5.5, .5.5.5.5, .5.5.5.5 ${\color[rgb]{0,0,1}-3.50\%}$ ${\color[rgb]{0,0,1}+0.53\%}$ ${\color[rgb]{0,0,1}+7.37\%}$ ${\color[rgb]{0,0,1}-5.38\%}$ ${\color[rgb]{1,0,0}-0.71\%}$ ${\color[rgb]{1,0,0}-2.30\%}$
$1$, $1$, 00 ${\color[rgb]{0,0,1}-2.52\%}$ ${\color[rgb]{0,0,1}+0.48\%}$ ${\color[rgb]{0,0,1}+6.07\%}$ ${\color[rgb]{0,0,1}-4.77\%}$ ${\color[rgb]{0,0,1}+6.03\%}$ ${\color[rgb]{1,0,0}-1.97\%}$
temperature .5.5.5.5, $1$, 00 ${\color[rgb]{0,0,1}-3.06\%}$ ${\color[rgb]{0,0,1}+0.19\%}$ ${\color[rgb]{0,0,1}+19.77\%}$ ${\color[rgb]{0,0,1}-6.86\%}$ ${\color[rgb]{1,0,0}-13.09\%}$ ${\color[rgb]{1,0,0}-2.76\%}$
.5.5.5.5, .5.5.5.5, .5.5.5.5 ${\color[rgb]{0,0,1}-2.37\%}$ ${\color[rgb]{0,0,1}+0.27\%}$ ${\color[rgb]{0,0,1}+11.79\%}$ ${\color[rgb]{0,0,1}-6.43\%}$ ${\color[rgb]{0,0,1}+13.72\%}$ ${\color[rgb]{1,0,0}-0.98\%}$
$1$, $1$, 00 ${\color[rgb]{0,0,1}-1.49\%}$ ${\color[rgb]{0,0,1}+0.27\%}$ ${\color[rgb]{0,0,1}+11.16\%}$ ${\color[rgb]{0,0,1}-7.11\%}$ ${\color[rgb]{1,0,0}-10.49\%}$ ${\color[rgb]{1,0,0}-2.64\%}$
@@
Table 3: Percentage of change in data quality when using our proposed mitigation strategy versus the baseline. Results are shown for top-$k$ decoding and sampling/temperature for different values of $\alpha$, $\beta$, and $\gamma$ (blue / red: positive / negative results, $\uparrow$ / $\downarrow$: higher / lower is better).
5.3 Results on collapse prevention
As explained in Section 4, we assess our approach by adopting the decoding strategy that caused the most significant model collapse, i.e. pure sampling for GPT-2 and temperature sampling for SmolLM2, and for a more direct comparison, we also conduct experiments using top-$k$ decoding. At each generation $i$, we compare against the baseline of training on all samples in the pool of data $\mathcal{D}^{i}$ (Eq. 3). We also provide an “Oracle” performance which represents a perfect machine-generated text detector that filters all synthetic samples.
We evaluate recursive training in the partially synthetic setting under $3$ mixing coefficient settings ($\alpha\!=\!1$, $\beta\!=\!1$, $\gamma\!=\!0$, $\alpha\!=\!0.5$, $\beta\!=\!1$, $\gamma\!=\!0$ and $\alpha\!=\!\beta\!=\!\gamma\!=\!0.5$). The task performance, data quality, and detector accuracy metrics over $9$ generations of partially synthetic training are depicted in Figure 2 for the scenario when there is an equal proportion of human and synthetic samples (we also provide results for the $2$ additional scenarios in Appendix S2,S3 ). Weighted sampling (as described in section 5.2) appears to prevent model collapse, while the baseline degrades (higher perplexity, lower test accuracy). Our method also alleviates any effects on the readability and diversity of the synthetic samples. The detectability of the baseline text, on the other hand, is increasing as it becomes more dissimilar to human text. Interestingly, our method improves performance compared to training exclusively on human data (Oracle) for both language models and respective decoding strategies. These outcomes demonstrate both the value of using synthetic data in LLM training, but also the importance of selecting the right synthetic samples while maintaining a high proportion of human data.
Figure 3: Distribution of perplexity for machine-generated text of GPT-2 at generations 00, $1$, and $9$ in the partially synthetic scenario $\left(\alpha\!=\!0.5,\beta\!=\!1,\gamma\!=\!0\right)$. Perplexity is evaluated using the model trained on the human text ($p^{0}$).
@@ 5.3 Results on collapse prevention
Table 3 enumerates the percentage difference for the final models under the baseline strategy vs. using our approach. We also enumerate the exact scores in Table S1. Our method improves the model output quality across all metrics compared to the baseline except for sampling with GPT-2. In this case, the diversity of the text increases dramatically using the baseline method, however, this is accompanied by a significant drop in MAUVE and readability compared to our method. We observe that mixing cross-generational data has an insignificant effect on the extent of model collapse compared to training solely on the previous generation, contrasting with the findings of Gerstgrasser et al. (2024). However, we note that Gerstgrasser et al. (2024) did not constrain the sample size and trained on all data available. Our experiments adopt the more realistic and less extreme setting, sampling the pool of data under mixing scenarios.
@@
Similar to Shumailov et al. (2024), we also study the perplexity distribution of the synthetic data at each generation using model $p^{0}$ that was trained on human data. Figure 3 depicts these distributions for generations 00, $1$, and $9$ for the GPT-2 model with top-$k$ and pure sampling, compared to the baseline. For top-$k$ decoding, similarly to Shumailov et al. (2024), we observe that for the baseline perplexity shifts towards regions of lower scores and the distribution becomes more peaked, showing signs of early model collapse. For pure sampling, on the other hand, we observe that the baseline distribution shifts to higher perplexity scores and becomes more spread out. This is an interesting finding that demonstrates that by removing truncation from the decoding strategy, the narrowing effect of model collapse is diminished and instead, model collapse is reflected by long-tail incoherent text that is completely distinct from the original human samples. By deploying our mitigation strategy, however, we observe very little change in the perplexity distribution for both sampling strategies, and also improved data quality metrics to the baseline, as seen in Table 3.

### v5

@@ 5 The impact of decoding strategies on model collapse
We carry out recursive training as described in section 4.1 on the open-ended text generation task by fine-tuning LLMs on the WikiText-2 dataset Merity et al. (2016) with GPT-2 $124$M  Radford et al. (2019) and SmolLM2 $360$Allal et al. (2025). The Wikipedia articles are segmented into non-overlapping chunks of $512$ tokens, where the first $256$ are used as the context ($\mathbf{x}$), and the remaining $256$ as the continuation ($\hat{\mathbf{x}}$). We conduct full fine-tuning for $1$ epoch and, to avoid information leakage between generation and training, define cross-entropy loss only on the generated sequence of each sample Dohmatob et al. (2024). Additional details can be found in Appendix A.3. We evaluate a range of decoding strategies to assess the effect on model collapse: greedy decoding ($\tau=0$), $5$-way beam search, pure sampling ($\tau=1$), temperature ($\tau=0.9$), top-$k$ (with $k=50$), and nucleus sampling ($\eta=0.95$). The hyperparameter settings for these decoding strategies were based on recommendations from prior work Holtzman et al. (2020); Shumailov et al. (2024); Arias et al. (2025).
@@
Figure 1 depicts the perplexity and evaluation accuracy on the WikiText-2 test set for every model generation. Additionally, we obtain scores for the qualitative metrics using the outputs generated by the model (i.e. $\{\hat{\mathbf{x}}\}_{s=1}^{n}$ of $\mathcal{D}_{\text{S}}^{i}$ in Algorithm 1), and enumerate them in Table 1 for generations 00 and $9$. We observe that deterministic decoding leads to the most severe model collapse. While stochastic sampling methods exhibit linear degradation across generations, collapse accelerates under greedy decoding and beam search before plateauing in later generations. At generation 00, deterministic strategies yield significantly less fluent and more repetitive text, with MAUVE scores below $5\%$ and diversity scores less than $20\%$. The disparity in generation quality between deterministic and stochastic strategies in the open-ended text generation task has been explored in related literature Holtzman et al. (2020); Pillutla et al. (2021). Here, we demonstrate that this disparity compounds across recursive training, resulting in significantly higher perplexity at generation $9$. While deterministic methods are rarely used in open-ended generation, we included them for completeness and due to the choice of beam-search in the experiments of Shumailov et al. (2024), but exclude them in subsequent experiments due to their unrealistic collapse.
@@
Pure sampling impacted the models differently. For GPT-2, sampling directly from the probability distribution produces diverse and fluent text at generation 00, but training recursively on these outputs results in the worst test perplexity among stochastic methods ($58.64$) and generated text that has low similarity to human text (MAUVE $7.18$). In contrast, with SmolLM2, pure sampling yields the smallest decline in task performance and maintains the closest overall similarity to human references across all evaluated metrics.
[…]
@@
In our subsequent experiments on preventing model collapse, we seek to validate that our method can work in the most extreme scenario. For this reason, we evaluate the models using the worst-performing stochastic decoding method (pure sampling for GPT-2 and temperature sampling for SmolLM2). In addition, to facilitate direct comparisons, we also evaluate with top-$k$ decoding due to the consistent performance across models.
@@ 6 Preventing model collapse
So far, we have carried out recursive training in a setting where models are trained exclusively on the outputs of the previous generation without implicitly including any human-generated samples. We now turn our focus to the partially synthetic setting, a more realistic scenario where human data make up a portion of the training dataset and the synthetic data is a mix of the samples produced across generations. The training dataset for generation $i$, $\mathcal{D}^{i}$, consists of the aggregation of $3$ samples:
@@
We explore the following settings: (i) fully synthetic $\left(\alpha\!=\!0,\beta\!=\!1,\gamma\!=\!0\right)$, where training data consists entirely of synthetic samples from the previous generation, (ii) partially synthetic $\left(\alpha\!>\!0,\beta\!=\!1,\gamma\!=\!0\right)$, where the same proportion of human data is added to the training data at every generation, and (iii) partially synthetic with synthetic data accumulated across generations $\left(\alpha\!=\!0.5,\beta\!=\!0.5,\gamma\!=\!0.5\right)$ as proposed in Gerstgrasser et al. (2024). We evaluate our method in the partially synthetic setting and vary the mixing coefficients $\alpha$, $\beta$, and $\gamma$, however, our method does not assume access to the values of the mixing coefficients and hence the data distribution. To prevent model collapse when the origin of each training sample is unknown, we train a machine-generated detector that estimates the likelihood of text origin (section 6.1). We then use this information to conduct importance sampling (section 6.2) that ultimately mitigates model collapse.
@@
6.1 Machine-generated text detection performance
We trained and evaluated RoBERTa, DeBERTav3 and ModernBERT on the MAGE dataset Li et al. (2024), a machine-generated text detection benchmark based on documents from $10$ domains which have been used to generate text from $27$ LLMs. We adopt the preset training / validation / test splits ($80\%$/$10\%$/$10\%$). We also test on the more demanding out-of-distribution test set that contains human-curated text from $4$ unseen domains and machine-generated samples by an unseen LLM (GPT-4). Each model was fine-tuned for 5 epochs using a binary cross-entropy loss. More details on the model training can be found in Appendix B. Performance is enumerated in Table 2. ModernBERT [^3] Fine-tuned ModernBERT machine-generated text detector, huggingface.co/GeorgeDrayson/modernbert-ai-detection. yielded the best classification performance on both aforementioned test sets with an AUC of .986.986.986.986 and .943.943.943.943, respectively. This is comparable to the top-performing model evaluated by Li et al. (2024), Longformer, which achieved an in-distribution and out-of-distribution AUC of .99.99.99.99 and .94.94.94.94.
@@ 6.2 Informed sampling of training data
Given a dataset at generation $i$, $\mathcal{D}_{i}\sim g(x)$, composed of an unknown mixture of human and synthetic samples, our goal is to sample a training dataset from a target human data distribution $\mathcal{D}_{\text{H}}\sim h(x)$ to prevent model collapse. We consider that a language model has collapsed if the inclusion of synthetic samples in the training data results in degraded performance compared to training exclusively on human samples.
We use Sampling Importance Resampling (SIR) Rubin (1988), a method for approximately sampling from a target distribution $h(x)$ based on sampling with importance weights from a proposal distribution $g(x)$ using the normalised likelihood ratio $h(x)/g(x)$. As this ratio is intractable in our case, we instead employ a machine-generated text detector to assign each sample $\mathbf{x}_{i}$, $\forall i$, a predicted probability $q\!\left(\mathbf{x}_{i}\right)$ of being machine-generated, which we treat as an approximation for the likelihood ratio.
As the detector has been trained on an unbalanced dataset ($29\%$ human samples), the predictions are biased towards attributing text as machine-generated, reflected in the optimal classification threshold of $0.8674$ (00/$1$: human/machine-generated). To ameliorate this, we apply a bias term $b\!\geq\!1$ (see Appendix C) to the probabilities, followed by normalising the weights using
@@
where $w_{i}\!\in\![0,1]$ denotes the weight for sample $\mathbf{x}_{i}$, $\forall i$. From the $n$ weighted training samples, we draw $k\!\times\!n$ samples with replacement, with $k=1.5$ to allow for a $50\%$ upsampling of the training data. In this way, we obtain a revised set of samples that we use in our recursive training regime.
6.3 Results on collapse prevention
As explained in section 5, we assess our approach by adopting the decoding strategy that caused the most significant model collapse, i.e. pure sampling for GPT-2 and temperature sampling for SmolLM2, and for a more direct comparison, we also conduct experiments using top-$k$ decoding. At each generation $i$, we compare against the baseline of training on all samples in the pool of data $\mathcal{D}^{i}$ (Eq. 3). We also provide an “Oracle” performance, which represents a perfect machine-generated text detector that filters all synthetic samples.
We evaluate recursive training in the partially synthetic setting under $3$ mixing settings ($\alpha\!=\!1$, $\beta\!=\!1$, $\gamma\!=\!0$), ($\alpha\!=\!0.5$, $\beta\!=\!1$, $\gamma\!=\!0$), ($\alpha\!=\!0.5$, $\beta\!=\!0.5$, $\gamma\!=\!0.5$). The task performance, data quality, and detector accuracy metrics over $10$ generations of partially synthetic training are depicted in Figure 2 for the scenario where human and synthetic samples have equal proportion (Appendices S4 and S6 contain results for the other $2$ scenarios). Weighted sampling (section 6.2) prevents model collapse and preserves the readability and diversity of the synthetic outputs, while the baseline degrades in task performance and data quality. The baseline generations become increasingly detectable, indicating divergence from human text. Notably, our method improves performance compared to training exclusively on human data (Oracle) across models and decoding strategies. These outcomes demonstrate both the value of using synthetic data in LLM training, but also the importance of selecting the right synthetic samples.
@@ 6.3 Results on collapse prevention
pure sampling .5.5.5.5, $1$, 00 ${\color[rgb]{0,0,1}-7.41\%}$ ${\color[rgb]{0,0,1}+1.50\%}$ ${\color[rgb]{1,0,0}-1.30\%}$ ${\color[rgb]{1,0,0}+36.71\%}$ ${\color[rgb]{0,0,1}+74.06\%}$ ${\color[rgb]{0,0,1}+50.23\%}$
@@
SmolLM2 $360$M top-$k$5.5.5.5, $1$, 00 -4.60% +0.68% +7.06% -1.17% +15.28% -0.73%
.5.5.5.5, .5.5.5.5, .$5$4.37% +0.62% +4.25% -0.42% +2.05% -0.74%
$1$, $1$, 00 -3.05% +0.54% +6.20% -1.74% +10.62% -0.61%
temperature .5.5.5.5, $1$, 00 -3.91% +0.55% +14.62% -1.46% +20.92% -1.86%
.5.5.5.5, .5.5.5.5, .$5$3.24% +0.48% +7.44% -1.98% -0.33% -1.35%
$1$, $1$, 00 -2.23% +0.39% +8.55% -1.80% +13.98% -0.34%
SmolLM2 $135$M top-$k$ $1$, $1$, 00 -4.19% +0.75% +9.96% -2.15% +13.08% -0.76%
SmolLM2 $1.7$B top-$k$ $1$, $1$, 00 -0.85% +0.06% +3.44% -1.11% +8.46% -0.65%
@@
Table 3: Percentage of change in data quality when using our proposed mitigation strategy versus the baseline. Results are shown for top-$k$ decoding and pure sampling/temperature for different values of $\alpha$, $\beta$, and $\gamma$ (blue / red: positive / negative results, $\uparrow$ / $\downarrow$: higher / lower is better).
@@
Table 3 enumerates the percentage difference for the final models under the baseline strategy vs. using our approach. We also enumerate the exact scores in Table S2. Our method improves the data quality and model performance across all metrics, except for pure sampling with GPT-2, where the baseline shows higher diversity and Self-BLEU but at the cost of lower MAUVE, readability, and task performance, indicating degraded quality. Notably, we observe that mixing cross-generational data has a minimal effect on the extent of model collapse compared to training solely on the previous generation, contrasting with the findings of Gerstgrasser et al. (2024). However, we note that Gerstgrasser et al. (2024) did not constrain the sample size. Our experiments adopt a more realistic and less extreme setting by sampling a fixed dataset under different mixing scenarios.
Additionally, we evaluate the effectiveness of our method at different model scales of SmolLM2 ($135$M, $360$M, $1.7$B) for a fixed decoding strategy and data mixing setting. The results are enumerated at the bottom of Table 3 and the performance across generations is depicted in Figure S5. Our method prevents model collapse across all SmolLM variants. Notably, we find that smaller models exhibit greater relative improvements using our method compared to both the Oracle and the baseline.
Figure 3: Perplexity distribution for machine-generated text of GPT-2 at generations 00, $1$, and $9$ in the partially synthetic scenario $\left(\alpha\!=\!0.5,\beta\!=\!1,\gamma\!=\!0\right)$. Perplexity is evaluated using the model trained on human text ($p^{0}$).
@@
In line with previous research Shumailov et al. (2024), we also study the perplexity distribution of the synthetic data at each generation using model $p^{0}$ that was trained on human data. Figure 3 depicts these distributions for generations 00, $1$, and $9$ for the GPT-2 model with top-$k$ and pure sampling, compared to the baseline (Appendix S3 contains the results for SmolLM2). For top-$k$ decoding, similarly to Shumailov et al. (2024), we observe that for the baseline, perplexity shifts towards regions of lower scores and the distribution becomes more peaked, showing signs of early model collapse. For pure sampling, on the other hand, we observe that the baseline distribution shifts to higher perplexity scores and displays increased variance. This is an interesting finding that demonstrates that by removing truncation from the decoding strategy, the narrowing effect of model collapse is diminished, and instead, model collapse is reflected by long-tail incoherent text that is completely distinct from the original human samples. By deploying our mitigation strategy, however, we observe very little change in the perplexity distribution for both sampling strategies.